# VLegal-Bench: Cognitively Grounded Benchmark for Vietnamese Legal Reasoning of Large Language Models

### v1

@@ 4.2. Evaluated Model
We evaluate VLegal-Bench using a total of 24 Large Language Models (LLMs), which are categorized into two main groups: general multilingual LLMs and domain-adapted Vietnamese LLMs. The general multilingual LLMs include both closed-source APIs, namely GPT-4o, GPT-4o-mini, Claude 4.5 Haiku, Claude 4.5 Sonnet, Gemini 2.5 Flash, and Gemini 2.5 Pro, as well as open-source models, including Qwen 2.5 Instruct (3B/7B/14B/32B/72B), Qwen 2.5 (3B/7B), Llama 2 Chat (7B/13B), Meta Llama 3 Instruct (8B/70B), InternLM 3 8B Instruct, and Gemma 2 Instruct (9B/27B). The Vietnamese-focused models include SeaLLMs v3 Chat (1.5B/7B), BloomVN 8B Chat, and a Vietnamese legal domain-adapted model, VLSP2025-LegalSML 4B. We provide detailed information about these models in Appendix B.
@@
Domain-adapted Vietnamese LLMs BloomVN-8B-chat 46.66 65.59 63.67 65.29 26.75 70.36 45.00 0.500 49.08 57.00
VLSP2025-qwen3-4b-legal-pretrain 62.83 75.99 66.67 70.35 24.50 79.45 45.00 0.716 73.29 58.69
@@
Domain-adapted Vietnamese LLMs BloomVN-8B-chat 32.17 82.33 70.89 0.00/39.61 50.00 0.2407 0.3188 0.4099 20.10 47.22 86.61 58.55
qwen3-4b-legal-pretrain 43.83 82.00 76.37 0.00/39.61 54.47 0.4361 0.3737 0.4381 34.54 61.57 89.08 59.40
@@ 4.3. Overall Performance Analysis
Even more striking is the performance of qwen3-4b-legal-pretrain, a 4-billion parameter model that achieves best-in-class results on article prediction (43.83%) and document summarization (0.4361 ROUGE-L), outperforming models up to 18 times its size. This demonstrates that targeted legal domain pretraining provides greater benefits than raw parameter scaling for specialized legal tasks.
@@ 4.6. Model Family Analysis
The three domain-adapted Vietnamese models show increasing returns with specialization. BloomVN-8B-chat, a general Vietnamese model, achieves moderate performance comparable to similarly-sized general multilingual models. VLSP2025-qwen3-4b-legal-pretrain demonstrates the power of legal domain pretraining, achieving best-in-class performance on article prediction and summarization despite having only 4 billion parameters. CMC-AI-Legal-32B represents the most specialized model, with legal-specific fine-tuning that yields dominant performance across reasoning and inference tasks (Level 3), achieving state-of-the-art results on four out of five tasks. This progression general Vietnamese → legal-pretrained → legal-fine-tuned validates a staged approach to building effective legal AI for low-resource languages.
@@ Appendix B Details of Tested LLMs
Model Type Model Size Context Len Access URL General Multilingual LLMs GPT-4o N/A 128k API https://platform.openai.com GPT-4o-mini N/A 128k API https://platform.openai.com Claude 2.5 Haiku N/A 200k API https://platform.claude.com/ Claude 2.5 Sonnet N/A 200k API https://platform.claude.com/ Gemini 2.5 Flash N/A 1M API https://ai.google.dev Gemini 2.5 Pro N/A 1M API https://ai.google.dev Qwen 2.5 Instruct 3B 32k Weights https://huggingface.co/Qwen Qwen 2.5 Instruct 7B / 14B / 32B / 72B 128k Weights https://huggingface.co/Qwen Qwen 2.5 3B 32k Weights https://huggingface.co/Qwen Qwen 2.5 7B 128k Weights https://huggingface.co/Qwen Llama 2 Chat 7B / 13B 4k Weights https://huggingface.co/meta-llama Meta Llama 3 Instruct 8B / 70B 8k Weights https://huggingface.co/meta-llama InternLM 3 Instruct 8B 8k Weights https://huggingface.co/internlm Gemma 2 Instruct 9B / 27B 8k Weights https://huggingface.co/google Vietnamese-focused LLMs SeaLLMs v3 Chat 1.5B 8k Weights https://huggingface.co/SeaLLMs SeaLLMs v3 Chat 7B 8k Weights https://huggingface.co/SeaLLMs BloomVN Chat 8B N/A Weights https://huggingface.co/BlossomsAI VLSP2025-LegalSML 4B 40960 Weights https://huggingface.co/VLSP2025-LegalSML

### v2

@@ 4.2. Evaluated Model
We evaluate VLegal-Bench using a total of 24 Large Language Models (LLMs), which are categorized into two main groups: general multilingual LLMs and domain-adapted Vietnamese LLMs. The general multilingual LLMs include both closed-source APIs, namely GPT-4o, GPT-4o-mini, Claude 4.5 Haiku, Claude 4.5 Sonnet, Gemini 2.5 Flash, as well as open-source models, including Qwen 2.5 Instruct (3B/7B/14B/32B/72B), Qwen 2.5 (3B/7B), Llama 2 Chat (7B/13B), Meta Llama 3 Instruct (8B/70B), InternLM 3 8B Instruct, and Gemma 2 Instruct (9B/27B). The Vietnamese-focused models include SeaLLMs v3 Chat (1.5B/7B), BloomVN 8B Chat, and a Vietnamese legal domain-adapted model CMC-AI-Legal-32B. We provide detailed information about these models in Appendix B.
@@
Domain-adapted Vietnamese LLMs SeaLLMs-v3-7B-Chat 62.23 68.96 57.67 58.88 22.50 76.28 49.67 0.475 62.10 54.39
BloomVN-8B-chat 46.66 65.59 63.67 65.29 26.75 70.36 45.00 0.500 49.08 57.00
@@
Domain-adapted Vietnamese LLMs SeaLLMs-v3-7B-Chat 26.00 81.50 64.38 0.00/39.61 49.16 0.1700 0.3547 0.4141 39.76 56.94 92.53 63.25
BloomVN-8B-chat 32.17 82.33 70.89 0.00/39.61 50.00 0.2407 0.3188 0.4099 20.10 47.22 86.61 58.55
@@ 4.6. Model Family Analysis
The three domain-adapted Vietnamese models show increasing returns with specialization. BloomVN-8B-chat, a general Vietnamese model, achieves moderate performance comparable to similarly-sized general multilingual models. CMC-AI-Legal-32B represents the most specialized model, with legal-specific fine-tuning that yields dominant performance across reasoning and inference tasks (Level 3), achieving state-of-the-art results on four out of five tasks. This progression general Vietnamese → legal-pretrained → legal-fine-tuned validates a staged approach to building effective legal AI for low-resource languages.
@@ Appendix B Details of Tested LLMs
Model Type Model Size Context Len Access URL General Multilingual LLMs GPT-4o N/A 128k API https://platform.openai.com GPT-4o-mini N/A 128k API https://platform.openai.com Claude 2.5 Haiku N/A 200k API https://platform.claude.com/ Claude 2.5 Sonnet N/A 200k API https://platform.claude.com/ Gemini 2.5 Flash N/A 1M API https://ai.google.dev Gemini 2.5 Pro N/A 1M API https://ai.google.dev Qwen 2.5 Instruct 3B 32k Weights https://huggingface.co/Qwen Qwen 2.5 Instruct 7B / 14B / 32B / 72B 128k Weights https://huggingface.co/Qwen Qwen 2.5 3B 32k Weights https://huggingface.co/Qwen Qwen 2.5 7B 128k Weights https://huggingface.co/Qwen Llama 2 Chat 7B / 13B 4k Weights https://huggingface.co/meta-llama Meta Llama 3 Instruct 8B / 70B 8k Weights https://huggingface.co/meta-llama InternLM 3 Instruct 8B 8k Weights https://huggingface.co/internlm Gemma 2 Instruct 9B / 27B 8k Weights https://huggingface.co/google Vietnamese-focused LLMs SeaLLMs v3 Chat 1.5B 8k Weights https://huggingface.co/SeaLLMs SeaLLMs v3 Chat 7B 8k Weights https://huggingface.co/SeaLLMs BloomVN Chat 8B 8k Weights https://huggingface.co/BlossomsAI